# Microsoft Industry AI

### v1

@@ Abstract
The application of large language models (LLMs) in domain-specific contexts, including finance, has expanded rapidly. Domain-specific LLMs are typically evaluated based on their performance in various downstream tasks relevant to the domain. In this work, we present a detailed analysis of fine-tuning LLMs for such tasks. Somewhat counterintuitively, we find that in domain-specific cases, fine-tuning exclusively on the target task is not always the most effective strategy. Instead, multi-task fine-tuning - where models are trained on a cocktail of related tasks - can significantly enhance performance. We demonstrate how this approach enables a small model, such as Phi-3-Mini, to achieve state-of-the-art results, even surpassing the much larger GPT-4-o model on financial benchmarks. Our study involves a large-scale experiment, training over 200 models using several widely adopted LLMs as baselines, and empirically confirms the benefits of multi-task fine-tuning. Additionally, we explore the use of general instruction data as a form of regularization, suggesting that it helps minimize performance degradation. We also investigate the inclusion of mathematical data, finding improvements in numerical reasoning that transfer effectively to financial tasks. Finally, we note that while fine-tuning for downstream tasks leads to targeted improvements in task performance, it does not necessarily result in broader gains in domain knowledge or complex domain reasoning abilities.
@@ 1 Introduction
In this work, we investigated the impact of multi-task fine-tuning. Instead of fine-tuning the model solely on the target downstream task, we fine-tune it on multiple related downstream tasks simultaneously. We conduct a massive ablation study to explore the interactions between various financial tasks and datasets. In total, we train 220 models to provide an in-depth evaluation of different financial benchmarks and LLMs. Our empirical findings demonstrate that incorporating training data from multiple downstream tasks creates a cocktail effect, where the integration of multiple datasets creates a synergistic improvement in model performance, even for a single task.
@@ 2.1 Background
Multi-task training is not a new concept. The efficiency of this approach has been demonstrated for general domains in the past (Aribandi et al., 2021; Aghajanyan et al., 2021). More recent work has shown success with instruction tuning specifically (Wang et al., 2023b; Yue et al., 2023), as well as showing the impact of additional datasets. On the other hand, the exact interactions between tasks are still understudied, especially in the domain-specific case, and more specifically for finance. Past approaches to domain-specific adaptation, such as Cheng et al. (2023), used broader domain data, removing the ability to observe the interactions between the tasks themselves. While Wang et al. (2023a) use a task oriented approach in finance, there is no measurement on the task level, or experimentation around adding general data.
@@ 3.1 Core Financial Datasets
  • Headline: This dataset consists of financial news headlines, accompanied by binary questions. The dataset aims to represent how financial information is presented in news media, and the primary purpose of the dataset is event detection in finance. This dataset is an adaptation of the original headline dataset (Sinha & Khandait, 2021) by FinGPT (Wang et al., 2023a).
@@ 4.1 Experiment Setup
All experiments were conducted using a single machine with 2 Nvidia H100 GPUs. All experiments were done using full fine-tuning of all weights in the model. We experimented with various learning rates, ranging from $3\text{e}^{-6}$ to $3\text{e}^{-5}$. We used three epochs for the smaller runs ($\binom{n}{1},\ \binom{n}{2}$), and two epochs for the rest. The longest single fine-tuning experiment took under three hours to run.
@@ 4.2 Metrics
[…]
@@
Numerical Evaluations: FinQA and ConvFinQA require evaluating numerical exact match (EM) for scoring. To prevent issues stemming from rounding errors, or scale representations, we used a heuristic relaxation of exact match. We say that $x$ is numerically same to $y$ if for some small $\epsilon$, $y\pm\epsilon=x^{n},n\in\{10^{-6},10^{-3},10^{-2},10^{0},10^{2},10^{3},10^{6}\}$. While not exhaustive, these are very common scales in finance (millions vs thousands vs billions, dollars vs cents, basis points, etc.).
Classification: To evaluate classification tasks we use standard accuracy scores.
@@
Table 2: Full experiment results for single-task and multi-task fine-tuning, aggregated across all experiments for three LLMs. Baseline results from the original models are provided for reference. The multi-task fine-tuning result represents the best performance across multi-task combinations.
Phi-3-Small Mistral-7B-Instruct-v0.3 Llama-3.1-8B-Instruct Baseline Single-task Multi-task Baseline Single-task Multi-task Baseline Single-task Multi-task Headline 0.67 0.67 0.96 0.69 0.67 0.95 0.53 0.67 0.95 FPB 0.48 0.86 0.89 0.78 0.67 0.89 0.76 0.82 0.89 FinNerCLS 0.71 0.96 0.98 0.66 0.97 0.98 0.54 0.97 0.99 FinQA 0.47 0.44 0.64 0.46 0.39 0.56 0.66 0.61 0.65 ConvFinQA 0.65 0.73 0.81 0.70 0.72 0.81 0.77 0.83 0.85 TwitterTopics 0.41 0.87 0.88 0.48 0.85 0.88 0.52 0.86 0.87 Twitter SA 0.65 0.85 0.91 0.80 0.83 0.91 0.68 0.80 0.91
@@ 4.3 Main Results
The Cocktail Effect: In Table 2, we present a comparison for the three LLMs using the metrics discussed above. A visualization of these results is provided in Fig. 3. It is clear that fine-tuning, whether single-task or multi-task, significantly improves performance compared to the baseline. Both fine-tuning approaches outperform the baseline across the vast majority of benchmarks, with this trend holding consistently across all three models.
@@
When aligning LLMs, Ouyang et al. (2022) adapt the loss use by Stiennon et al. (2020), including a regularization term: $\beta\log\left[\mathcal{M}_{\text{RL},\phi}(y|x)/\mathcal{M}_{\text{SFT}}(y|x)\right]$. This component is used to ensure the new model does not stray ’too far’ from the original model, and is missing in the standard domain adaptation regime. We hypothesize that since the pretrained model $\mathcal{M}$ has already been exposed to Open-Orca, incorporating it in finetuning serves a similar purpose. In other words, we assume:
$\log\left[\mathcal{M}_{\mathcal{D}_{\text{domain}}}(y|x)/{\mathcal{M}(y|x)}\right]\geq\log\left[\mathcal{M}_{(\mathcal{D}_{\text{domain}}\cup\mathcal{D}_{\text{gen}})}(y|x)/\mathcal{M}(y|x)\right].$
@@ 5 Related Work
Domain-specific LLMs: Recent advances in LLMs have led to many attempts at creating models tailored to specific domains. These models aim to outperform general-purpose ones by having deeper knowledge of the domain, being more effective at solving tasks relevant to that domain, or adopting a more appropriate style. Several methods have been suggested for training these models. One approach is to pre-train a language model entirely on domain-specific data, as seen in (Wu et al., 2023; Singhal et al., 2023). Another common approach is to take pre-trained LLMs and fine-tune them for specific downstream tasks (Xie et al., 2023; Wang et al., 2023a; Cheng et al., 2024; Jiang et al., 2024; Cheng et al., 2023) in a domain adaptation process.
Domain Adaptation of LLMs: Various techniques have been developed to transform a general language model into a domain-specific one. One option is continual pre-training (CPT) (Gururangan et al., 2020), where a pre-trained LLM undergoes further training on raw data that contains relevant domain-specific knowledge, enhancing the model’s understanding of that domain. Another method involves supervised fine-tuning (SFT), where the model is trained on a large set of domain-specific instructions (Wei et al., 2021). Some approaches focus on specific tasks within the domain, fine-tuning the model with instruction datasets tailored to those particular tasks (Wang et al., 2023a). Additionally, a hybrid approach has been proposed, where CPT is performed first, followed by domain-specific instruction tuning to refine the model’s capabilities (Bhatia et al., 2024; Wu et al., 2024; Xie et al., 2024b).
@@ 6 Conclusions
In this work, we demonstrated the potential of multi-task fine-tuning as a robust approach to optimizing the performance of LLMs on downstream tasks. Through extensive experimentation involving over 200 models, we showed that combining training data from multiple related financial tasks creates a ”cocktail effect”, yielding significant performance gains, and even allowing smaller models such as Phi-3-Mini to surpass larger counterparts like GPT-4-o on targeted benchmarks. Our findings highlight the advantages of a training approach that leverages synergies between tasks.
@@ Appendix A LLM as a Judge Prompt
For example, if the reference answer is ”It cost \$5B annually” and the assistant answer is ”It cost \$5 billion per year”, the rating should be 2.
If the assistant answer is ”It cost \$5”, the rating should be 1.
If the assistant answer is ”It cost \$4 million per month”, the rating should be 0.
@@
For example, if the question is ”What was his salary?” and the reference answer is ”We can see that by adding the various components in table 3, we get that 3K + 7.5K equals a total salary of 10.5K annually”, and the assistant’s answer is ”10,500”, the rating should be 2.
If the assistant’s answer is ”10.5K. This salary reflects and excellent compensation given the low cost of living in the area”, the rating should still be 2.
If the assistant’s answer is ”the answer can be found in table 3 by adding 3K + 7.5K”, the rating should be 1.
If the assistant’s answer is ”7.5K”, the rating should be 0.
@@ Appendix E Dataset Examples
This LOAN AND SECURITY AGREEMENT dated January 27, 1999, between SILICON VALLEY BANK (”Bank”), a California-chartered bank with its principal place of business at 3003 Tasman Drive, Santa Clara, California 95054 with a loan production office located at 40 William St., Ste.
@@
You are given a financial sentence taken from Twitter. Your task is to infer its topic. Answer using one of the following labels: [’Analyst Update’, ’Fed — Central Banks’, ’Company — Product News’, ’Treasuries — Corporate Debt’, ’Dividend’, ’Stock Commentary’, ’Stock Movement’], and include nothing else.
@@
The following are multiple choice questions (with answers) about economics. Think step by step and then finish your answer with ”the answer is (X)” where X is the correct letter choice.
@@
Mr. Jones is president of the First National Bank of St. Louis and wishes to determine if his bank is holding too much of its demand deposits as reserves. The bank’s total deposits = \$1,700,000 and the reserve ratio is 20%. If Mr. Jones finds that reserves = \$850,000 what might he conclude about excess reserves? Options: A: ”\$340,000”, B: ”\$600,000”, C: ”\$425,000”, D: ”25%”, E: ”10%”, F: ”\$510,000”, G: ”\$1,700,000”, H: ”30%”, I: ”\$255,000”, J: ”15%”
@@
The following are multiple choice questions (with answers) about business. Think step by step and then finish your answer with ”the answer is (X)” where X is the correct letter choice.
@@
Mr. Frankel wants to borrow \$2,000 from November 16 for 143 days. The interest rate is 6%. What would the difference in the interest charge amount to if the bank used exact interest instead of bankers’ interest? Options: A: ”\$2.00”, B: ”\$0.25”, C: ”\$1.50”, D: ”\$1.32”, E: ”\$3.30”, F: ”\$0.50”, G: ”\$0.99”, H: ”\$0.66”, I: ”\$1.98”, J: ”\$2.64”

### v2

@@ Abstract
The application of large language models (LLMs) in domain-specific contexts, including finance, has expanded rapidly. Domain-specific LLMs are typically evaluated based on their performance in various downstream tasks relevant to the domain. In this work, we present a detailed analysis of fine-tuning LLMs for such tasks. Somewhat counterintuitively, we find that in domain-specific cases, fine-tuning exclusively on the target task is not always the most effective strategy. Instead, multi-task fine-tuning - where models are trained on a cocktail of related tasks - can significantly enhance performance. We demonstrate how this approach enables a small model, such as Phi-3-Mini, to achieve state-of-the-art results, even surpassing the much larger GPT-4-o model on financial benchmarks. Our study involves a large-scale experiment, conducting over 200 training experiments using several widely adopted LLMs as baselines, and empirically confirms the benefits of multi-task fine-tuning. Additionally, we explore the use of general instruction data as a form of regularization, suggesting that it helps minimize performance degradation. We also investigate the inclusion of mathematical data, finding improvements in numerical reasoning that transfer effectively to financial tasks. Finally, we note that while fine-tuning for downstream tasks leads to targeted improvements in task performance, it does not necessarily result in broader gains in domain knowledge or complex domain reasoning abilities.
Mixing It Up:
The Cocktail Effect of Multi-Task Fine-Tuning on LLM Performance - A Case Study in Finance
Meni Briefthanks: Equal contribution, corresponding authors: {menibrief, odedovadia}@microsoft.com       Oded Ovadia11footnotemark: 1
Gil Shenderovitz       Noga Ben Yoash       Rachel Lemberg       Eitam Sheetrit Microsoft Industry AI
@@ 1 Introduction
In this work, we investigated the impact of multi-task fine-tuning. Instead of fine-tuning the model solely on the target downstream task, we fine-tune it on multiple related downstream tasks simultaneously. We conduct a massive ablation study to explore the interactions between various financial tasks and datasets. In total, we conduct 220 training experiments to provide an in-depth evaluation of different financial benchmarks and LLMs. Our empirical findings demonstrate that incorporating training data from multiple downstream tasks creates a cocktail effect, where the integration of multiple datasets creates a synergistic improvement in model performance, even for a single task.
@@ 2.1 Background
Multi-task learning is not a new concept (Caruana, 1997). The efficiency of this approach has been demonstrated across various machine learning architectures in the past (Crawshaw, 2020). This is also true for general domains in natural language processing (Aribandi et al., 2021; Aghajanyan et al., 2021; Liu et al., 2019). More recent work has shown success with instruction tuning specifically (Wang et al., 2023b; Yue et al., 2023), as well as showing the impact of additional datasets. On the other hand, the exact interactions between tasks are still understudied, especially in the domain-specific case, and more specifically for finance. Past approaches to domain-specific adaptation, such as Cheng et al. (2023), used broader domain data, removing the ability to observe the interactions between the tasks themselves. While Wang et al. (2023a) use a task oriented approach in finance, there is no measurement on the task level, or experimentation around adding general data.
@@ 3.1 Core Financial Datasets
  • Headline: This dataset consists of financial news headlines, accompanied by binary questions. The dataset aims to represent how financial information is presented in news media, and the primary purpose of the dataset is event detection in finance. This dataset is an adaptation of the original headline dataset (Sinha and Khandait, 2021) by FinGPT (Wang et al., 2023a).
@@ 4.1 Experiment Setup
All experiments were conducted using a single machine with 2 Nvidia H100 GPUs. All experiments were done using full fine-tuning of all weights in the model. We experimented with various learning rates, ranging from $3\text{e}^{-6}$ to $3\text{e}^{-5}$. We used three epochs for the smaller runs ($\binom{n}{1},\ \binom{n}{2}$), and two epochs for the rest. The longest single fine-tuning experiment took under three hours to run. This choice of hyperparameters made sure that all training runs converged well, thus enabling a fair comparison. Following the process described in Section 2.2 and using the nine datasets listed in Section 3, we ended up with 55 unique training dataset mixes, resulting in 55 distinct training runs for each of the four models - yielding a total of 220 different experiments. We have made the code for running the training process available at https://anonymous.4open.science/r/cocktail_effect-54F8/README.md.
@@ 4.2 Metrics
$\begin{split}\text{Multi-task Score}&:=\max_{\mathcal{D}_{i}}\left(\mathcal{E}_{T_{i}}\left(\mathcal{M}_{\mathcal{D}_{i}}\right)\right)\\ &=\mathcal{E}_{T_{i}}\left(\mathcal{M}_{\mathcal{D}_{i}^{*}}\right)\end{split}$ (3)
@@
Numerical Evaluations: FinQA and ConvFinQA require evaluating numerical exact match (EM) for scoring. To prevent issues stemming from rounding errors, or scale representations, we used a heuristic relaxation of exact match. We say that $x$ is numerically same to $y$ if for some small $\epsilon$, it holds that $y\pm\epsilon=x^{n},n\in\{10^{-6},10^{-3},10^{-2},10^{0},10^{2},10^{3},10^{6}\}$. While not exhaustive, these are very common scales in finance (millions vs thousands vs billions, dollars vs cents, basis points, etc.).
Classification: To evaluate classification tasks we used standard (binary) accuracy scores.
@@
Table 2: Full experiment results for single-task and multi-task fine-tuning, aggregated across all experiments for three LLMs. Baseline results from the original models are provided for reference. The multi-task fine-tuning result represents the best performance across multi-task combinations. Margins of error are included for reference ($\alpha=0.01$).
Phi3-Small Mistral-7B-Instruct-v0.3 Llama-3.1-8B-Instruct Baseline Single-task Multi-task Baseline Single-task Multi-task Baseline Single-task Multi-task Headline 0.67$\pm$0.009 0.67$\pm$0.009 0.96$\pm$0.004 0.69$\pm$0.008 0.67$\pm$0.009 0.95$\pm$0.004 0.53$\pm$0.009 0.67$\pm$0.009 0.95$\pm$0.004 FPB 0.48$\pm$0.041 0.86$\pm$0.029 0.89$\pm$0.026 0.78$\pm$0.034 0.67$\pm$0.039 0.89$\pm$0.026 0.76$\pm$0.035 0.82$\pm$0.032 0.89$\pm$0.026 FinNerCLS 0.71$\pm$0.02 0.96$\pm$0.009 0.98$\pm$0.006 0.66$\pm$0.021 0.97$\pm$0.007 0.98$\pm$0.006 0.54$\pm$0.022 0.97$\pm$0.007 0.99$\pm$0.004 FinQA 0.47$\pm$0.038 0.44$\pm$0.038 0.53$\pm$0.038 0.46$\pm$0.038 0.39$\pm$0.038 0.47$\pm$0.038 0.66$\pm$0.036 0.61$\pm$0.038 0.62$\pm$0.037 ConvFinQA 0.65$\pm$0.032 0.73$\pm$0.03 0.81$\pm$0.026 0.70$\pm$0.031 0.72$\pm$0.03 0.81$\pm$0.026 0.77$\pm$0.028 0.83$\pm$0.025 0.85$\pm$0.024 TwitterTopics 0.41$\pm$0.02 0.87$\pm$0.014 0.88$\pm$0.013 0.48$\pm$0.02 0.85$\pm$0.014 0.88$\pm$0.013 0.52$\pm$0.02 0.86$\pm$0.014 0.87$\pm$0.014 Twitter SA 0.65$\pm$0.025 0.85$\pm$0.019 0.91$\pm$0.015 0.80$\pm$0.021 0.83$\pm$0.02 0.91$\pm$0.015 0.68$\pm$0.025 0.80$\pm$0.021 0.91$\pm$0.015
@@ 4.3 Main Results
The Cocktail Effect: In Table 2, we present a comparison for the three LLMs using the metrics discussed above. A visualization of these results is provided in Fig. 3. It is clear that fine-tuning, whether single-task or multi-task, significantly improves performance compared to the baseline. Both fine-tuning approaches outperform the baseline across the vast majority of benchmarks, with this trend holding consistently across all three models. Margins of error were calculated in the standard way, i.e. $z_{\frac{\alpha}{2}}\sqrt{\sigma^{2}/n}$.
@@
When aligning LLMs, Ouyang et al. (2022) adapt the loss used by Stiennon et al. (2020), including a regularization term: $\beta\log\left[\mathcal{M}_{\text{RL},\phi}(y|x)/\mathcal{M}_{\text{SFT}}(y|x)\right]$. This component is used to ensure the new model does not stray ’too far’ from the original model, and is missing in the standard domain adaptation regime. We hypothesize that since the pretrained model $\mathcal{M}$ has already been exposed to Open-Orca, incorporating it in finetuning serves a similar purpose. In other words, we assume:
$\log\left[\mathcal{M}_{\mathcal{D}_{\text{domain}}}(y|x)/{\mathcal{M}(y|x)}\right]\geq\\ \log\left[\mathcal{M}_{(\mathcal{D}_{\text{domain}}\cup\mathcal{D}_{\text{gen}})}(y|x)/\mathcal{M}(y|x)\right]$ (5)
@@ 5 Related Work
Domain-specific LLMs: Recent advances in LLMs have led to many attempts at creating models tailored to specific domains. These models aim to outperform general-purpose ones by having deeper knowledge of the domain, being more effective at solving tasks relevant to that domain, or adopting a more appropriate style. Several methods have been suggested for training these models. One approach is to pre-train a language model entirely on domain-specific data, as seen in (Wu et al., 2023; Singhal et al., 2023). Another common approach is to take pre-trained LLMs and fine-tune them for specific downstream tasks (Xie et al., 2023b; Wang et al., 2023a; Cheng et al., 2024; Jiang et al., 2024; Cheng et al., 2023) in a domain adaptation process.
Domain Adaptation of LLMs: Various techniques have been developed to transform a general language model into a domain-specific one. One option is continual pre-training (CPT) (Gururangan et al., 2020), where a pre-trained LLM undergoes further training on raw data that contains relevant domain-specific knowledge, enhancing the model’s understanding of that domain. Another method involves supervised fine-tuning (SFT), where the model is trained on a large set of domain-specific instructions (Wei et al., 2021). Some approaches focus on specific tasks within the domain, fine-tuning the model with instruction datasets tailored to those particular tasks (Wang et al., 2023a). There are also various works on approaches for selecting data for training (Xie et al., 2023a; Xia et al., 2024).Additionally, a hybrid approach has been proposed, where CPT is performed first, followed by domain-specific instruction tuning to refine the model’s capabilities (Bhatia et al., 2024; Wu et al., 2024; Xie et al., 2024b, c).
@@ 6 Conclusions
In this work, we demonstrated the potential of multi-task fine-tuning as a robust approach to optimizing the performance of LLMs on downstream tasks. Through extensive experimentation involving over 200 training runs, we showed that combining training data from multiple related financial tasks creates a "cocktail effect", yielding significant performance gains, and even allowing smaller models such as Phi-3-Mini to surpass larger counterparts like GPT-4-o on targeted benchmarks. Our findings highlight the advantages of a training approach that leverages synergies between tasks.
@@
Limitations
We acknowledge several limitations of this work. As with all experiments involving fine-tuning, the choice of hyperparameters plays a critical role. While we conducted a targeted hyperparameter search, the large scale of our experiments made a comprehensive grid search infeasible.
Additionally, the financial domain is vast, encompassing many intricacies and complexities that extend beyond the scope of the seven core datasets used in this study. Our work serves as a case study focusing on these representative datasets, but addressing other aspects of finance will necessitate the use of additional datasets tailored to those specific areas.
Finally, we note that while there are plenty of empirical results that demonstrate the general effectiveness of multi-task learning, there is still a significant lack of modern theory (Crawshaw, 2020). Although past works provide strong theoretical frameworks for multi-task learning (Evgeniou and Pontil, 2004; Ciliberto et al., 2015), it is difficult to extend them elegantly to modern deep learning methods.
@@ Appendix A LLM as a Judge Prompt
For example, if the reference answer is "It cost \$5B annually" and the assistant answer is "It cost \$5 billion per year", the rating should be 2.
If the assistant answer is "It cost \$5", the rating should be 1.
If the assistant answer is "It cost \$4 million per month", the rating should be 0.
@@
For example, if the question is "What was his salary?" and the reference answer is "We can see that by adding the various components in table 3, we get that 3K + 7.5K equals a total salary of 10.5K annually", and the assistant’s answer is "10,500", the rating should be 2.
If the assistant’s answer is "10.5K. This salary reflects and excellent compensation given the low cost of living in the area", the rating should still be 2.
If the assistant’s answer is "the answer can be found in table 3 by adding 3K + 7.5K", the rating should be 1.
If the assistant’s answer is "7.5K", the rating should be 0.
@@ Appendix C Full Results
Fig. 5 is a visualization of the results from Table 2, and shows the full results for each model across all seven tasks. Phi-3-Mini is brought here as well for completeness.
@@ Appendix E Dataset Examples
This LOAN AND SECURITY AGREEMENT dated January 27, 1999, between SILICON VALLEY BANK ("Bank"), a California-chartered bank with its principal place of business at 3003 Tasman Drive, Santa Clara, California 95054 with a loan production office located at 40 William St., Ste.
@@
You are given a financial sentence taken from Twitter. Your task is to infer its topic. Answer using one of the following labels: [’Analyst Update’, ’Fed - Central Banks’, ’Company - Product News’, ’Treasuries - Corporate Debt’, ’Dividend’, ’Stock Commentary’, ’Stock Movement’], and include nothing else.
@@
The following are multiple choice questions (with answers) about economics. Think step by step and then finish your answer with "the answer is (X)" where X is the correct letter choice.
@@
Mr. Jones is president of the First National Bank of St. Louis and wishes to determine if his bank is holding too much of its demand deposits as reserves. The bank’s total deposits = \$1,700,000 and the reserve ratio is 20%. If Mr. Jones finds that reserves = \$850,000 what might he conclude about excess reserves? Options: A: "\$340,000", B: "\$600,000", C: "\$425,000", D: "25%", E: "10%", F: "\$510,000", G: "\$1,700,000", H: "30%", I: "\$255,000", J: "15%"
@@
The following are multiple choice questions (with answers) about business. Think step by step and then finish your answer with "the answer is (X)" where X is the correct letter choice.
@@
Mr. Frankel wants to borrow \$2,000 from November 16 for 143 days. The interest rate is 6%. What would the difference in the interest charge amount to if the bank used exact interest instead of bankers’ interest? Options: A: "\$2.00", B: "\$0.25", C: "\$1.50", D: "\$1.32", E: "\$3.30", F: "\$0.50", G: "\$0.99", H: "\$0.66", I: "\$1.98", J: "\$2.64"